A piston- cylinder device contains 0.95 kg of oxygen initially at a Temperature of 27 0C and pressure due to the weight of 150 kPa. Heat is added to the gas until it reaches a temperature of 627 oC. Determine the amount of heat added during the process
1
Expert's answer
2021-07-25T09:17:19-0400
"Q=mc_p(T_2-T_1)\\Q=(0.95)(0.918)(627-27)=523\ kJ"